OpenStudy (anonymous):

Circumference of a Coin: The $1 coin here depicts Sacagawea and her infant son. The diameter of the coin is 26.5 millimeters. Find the circumference. Use 3.14 for π and round to at least 1 decimal place. The circumference is millimeters.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the requested sircumference, is given by the subsequent formula: \[\Large \begin{gathered} circumference = 2 \times \pi \times R = \hfill \\ = \frac{{2 \times 3.14 \times 26.5}}{2} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

L=2*π*r, where r=0.0265m, than L=2*π*0.0265m<=>L=π0.053m